OpenStudy (anonymous):

A number minus 20 times its reciprocal is 1. Find the number. If the problem has two solutions, write your answers in the box below.

OpenStudy (anonymous):

x = -4, x = 5

OpenStudy (asnaseer):

you need to turn the word problem into algebra first, so: for "A number" use, say, the symbol "x" "reciprocal" means one over the number 20 times just means 20 times the number so we get:\[x-20*\frac{1}{x}=1\]\[x-\frac{20}{x}=1\]multiply both sides by "x" to get:\[x^2-20=x\]\[x^2-x-20=0\]now factorise:\[(x-5)(x+4)=0\]so the answer is, as @robtobey stated, x=-4 or x=5

OpenStudy (cathyangs):

is this x-(20*1/x) or (x-20)*1/x? It makes a difference.

OpenStudy (asnaseer):

good question - it is worded a little ambiguously

OpenStudy (asnaseer):

but I guess from the simple answers we got it must be what we assumed it to be.

OpenStudy (anonymous):

x=5